A compound includes only Phosphorous and Sulpur.There are 1.24 grams of Phosphorous in 2.20 grams of the compound.What is the molecular formula of the compound?
Please give me a detailed answer
1
Expert's answer
2014-03-21T10:41:08-0400
Calculate the mass of sulfur in the compound:
m(S) = m(compound) - m(P) = 0.96 g
To find the molecular formula we should obtain the molar ratio between Sulfur and Phosphorus.
Find the amount of moles of Sulfur and Phosphorus in the compound:
n(S) = m/M = 0.96/32 = 0.03
n(P) = m/M = 1.24/31 = 0.04
n(S) : n(P) = 0.03 : 0.04 = 3 : 4
Hence the molecular formula of the compound is P4S3.